# Qini Curve Estimation Under Clustered Network Interference

### v1

@@ A.3 Expressing $q_{ij}$ in terms of the propensity score
A.4 Proof of unbiasedness of fractional IPW estimator
@@
A.5 Variance of additive IPW estimator
@@
$\widehat{\phi}^{\text{addIPW}}(\pi)=\frac{1}{N}\sum_{i=1}^{N}\left\{\sum_{j=1}^{M_{i}}\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{1}(X_{i})}-(M_{i}-1)\right\}Y_{i}~{}.$
As clusters are independent, we can write $\mathrm{Var}\left(\widehat{\phi}^{\text{addIPW}}\right)=\frac{1}{N^{2}}\sum_{i=1}^{N}\mathrm{Var}\left(\left\{\sum_{j=1}^{M_{i}}\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{1}(X_{i})}-(M_{i}-1)\right\}Y_{i}\right)$ where the variance terms inside the sum can be decomposed as
$\underbrace{\mathbb{E}\left[\left(\left\{\sum_{j=1}^{M_{i}}\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{1}(X_{i})}-(M_{i}-1)\right\}Y_{i}\right)^{2}\right]}_{(a)}-\underbrace{\mathbb{E}\left[\left\{\sum_{j=1}^{M_{i}}\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{1}(X_{i})}-(M_{i}-1)\right\}Y_{i}\right]^{2}}_{(b)}~{}.$
@@
$\mathrm{Var}\left(\widehat{\phi}^{\text{addIPW}}\right)=\frac{1}{N^{2}}\sum_{i=1}^{N}\left(\mathbb{E}\left[\left\{\sum_{j=1}^{M_{i}}\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{1}(X_{i})}-(M_{i}-1)\right\}^{2}Y_{i}^{2}\right]\right)-V(\pi)^{2}~{},$
we can see that the variance will scale quadratically with $M_{i}$.
Appendix B Estimating Qini curves in settings with no interference
@@
Note that the above assumption is a special case of assumption 5.1 with the exposure mapping $d_{ij}(\mathbf{W}_{i})=W_{ij}$.
@@
We can show that under assumption 3.1 and B.1, the above estimators are unbiased estimators for the policy value $V(\pi)$ and $C(\pi)$, respectively. Namely, we can show this with the same proof as in Appendix A.4, but replacing $d_{ij}(\mathbf{W_{i}})=[W_{ij},\overline{W}_{i}]$ with $d_{ij}(\mathbf{W_{i}})=W_{ij}$.
Appendix C Experimental details
C.1 Simulating marketplace dataset
@@
Next, to introduce heterogeneous treatment effects, we compute an item attractiveness score matrix $\mathbf{A}$ where element $A_{ij}$ relates buyer $i$’s interest in purchasing item $j$. This matrix depends on both the covariates and treatment as follows $A_{ij}=\delta_{ij}\cdot(A_{ij}^{(0)}+W_{ij}\cdot A_{ij}^{(1)})$, where $A_{ij}^{(w)}=X_{i}^{\top}\Omega_{w}Z_{ij}$ with $\Omega_{w}\sim\text{U}([0,1]^{12\times 11})$ for $w\in\{0,1\}$. The variable $d_{ij}\sim\text{Bern}(0.5)$ randomly masks some elements in $A_{ij}^{(1)}$ to zero; this emulates that some items will not respond at all to a treatment. Here, $A_{ij}$ typically lies in the range $[0,1]$, but if necessary we clip it to this range so that we later could interpret it as a probability.
C.2 Details on experiment shown in Figure 1
@@
We simulate a dataset with $N=20.000$ buyers (i.e., clusters) and $K=3$ items (i.e., units) per buyer with the exponential decay function. We estimate the Qini curve for a fixed treatment prioritization rule using the naive estimator and the IPW estimator. This was repeated 200 times and we plotted the average Qini curve.
C.3 Additional experimental results
In this section, we include additional experimental results, see Figure 4 and Table 1 to support our conclusions in Section 6 in the main paper.
@@ C.3 Additional experimental results
Figure 4: Comparing variance of each estimation strategy as we vary the number the buyers $N$ (i.e., clusters) or items $M$ (i.e., units) with $\eta_{\text{exp-decay}}$. While varying one, the other is kept fixed to either $N=20.000$ or $K=11$. The variance is reported over 200 repetitions.
Table 1: Comparison of ability to rank policies by each estimation strategy. We used $N=20.000$ number of buyers (i.e., clusters) and $K=11$ items (i.e., units) per buyer. We report the average Kendall rank correlation with respect to the ground truth ranking over 200 repetitions. Higher is better, where 1 corresponds to a perfect rank correlation.
Max Product Exponential decay
Estimation strategy
Naive (ignore interference) 1.000 0.808 1.000
Standard IPW 0.928 0.845 0.945
Fractional IPW 0.991 0.806 0.995
$\beta$-IPW ($\beta=1$) 1.000 1.000 1.000
$\beta$-IPW ($\beta=2$) 1.000 0.995 1.000

### v2

@@ A.3 Expressing $q_{ij}$ in terms of the propensity score
A.4 Proof of Theorem 2
@@
A.5 Proof of Lemma 2
Proof.
@@ A.5 Proof of Lemma 2
$\widehat{\phi}^{\text{addIPW}}(\pi)=\frac{1}{N}\sum_{i=1}^{N}\left\{\sum_{j=1}^{M_{i}}\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{1}(X_{i})}-(M_{i}-1)\right\}\tilde{Y}_{i}~.$
As clusters are independent, we can write $\mathrm{Var}\left(\widehat{\phi}^{\text{addIPW}}\right)=\frac{1}{N^{2}}\sum_{i=1}^{N}\mathrm{Var}\left(\left\{\sum_{j=1}^{M_{i}}\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{1}(X_{i})}-(M_{i}-1)\right\}\tilde{Y}_{i}\right)$ where the variance terms inside the sum can be decomposed as
$\underbrace{\mathbb{E}\left[\left(\left\{\sum_{j=1}^{M_{i}}\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{1}(X_{i})}-(M_{i}-1)\right\}\tilde{Y}_{i}\right)^{2}\right]}_{(a)}-\underbrace{\mathbb{E}\left[\left\{\sum_{j=1}^{M_{i}}\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{1}(X_{i})}-(M_{i}-1)\right\}\tilde{Y}_{i}\right]^{2}}_{(b)}~.$
@@
$\mathrm{Var}\left(\widehat{\phi}^{\text{addIPW}}\right)=\frac{1}{N^{2}}\sum_{i=1}^{N}\left(\mathbb{E}\left[\left\{\sum_{j=1}^{M_{i}}\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{1}(X_{i})}-(M_{i}-1)\right\}^{2}\tilde{Y}_{i}^{2}\right]\right)-V(\pi)^{2}~,$
we can see that the variance will scale quadratically with $M_{i}$. ∎
Appendix B Variance reduction through augmented weighted estimators
In this work, we have considered weighting estimators that require only a single nuisance model, namely the propensity score $e_{w}(x)=\Pr(W=w\mid X=x)$. Because the propensity score is assumed to be known in our setting, these estimators are appealing as they eliminate the need to estimate any additional nuisance models. This, in turn, helps avoid potential biases that could arise from model misspecification. However, a key limitation of weighting estimators is that they can suffer from high variance. In this appendix, we briefly discuss a technique that introduces an additional nuisance model without risking bias, while offering the potential for variance reduction if the added model is well specified. We also present experimental results showing that this approach can indeed reduce variance without introducing bias, although the magnitude of the reduction may vary depending on the setting.
We start with a weighting estimator of the form $\widehat{\phi}(\pi;\omega)=\frac{1}{N}\sum_{i=1}^{N}\omega(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})\tilde{Y}_{i}$, where $\omega$ is a pre-specified weighting function. Both the standard IPW estimator and the $\beta$-IPW estimator can be expressed in this form using different choices of weighting functions:
$\omega^{\text{IPW}}(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})=\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\quad\text{and}\quad\omega^{\beta\text{-IPW}}(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})=\sum_{\mathcal{U}\in\mathcal{I}_{i}^{\beta}}\prod_{j\in\mathcal{U}}\left(\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{\pi_{ij}}(X_{i})}-1\right)~.$ (11)
This class of weighting estimators does not include the fractional IPW estimator, since it performs weighting directly on unit-level outcomes, as can be seen from inspecting (9).
We shall focus on an “augmented” variant of these weighted estimators which we define as
$\widehat{\phi}^{\text{augmented}}(\pi;\omega)=\frac{1}{N}\sum_{i=1}^{N}\omega(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})(\tilde{Y}_{i}-\widehat{g}(X_{i}))+\widehat{g}(X_{i})~,$ (12)
where $\widehat{g}(x)$ is an estimator of the cluster-level conditional expectation of the outcome given the pre-treatment covariates, $\mathbb{E}[\tilde{Y}\mid X=x]$. The cluster-level treatment information is excluded to avoid modeling the potentially high-dimensional structure of the cluster-level treatment.
The augmented weighted estimator is reminiscent of the augmented inverse probability weighting (AIPW) estimator (Robins et al., 1994), which has been used to adjust for pre-treatment covariates with the goal of reducing estimator variance while protecting against bias when the true propensity score is known (see, e.g., Cao et al. (2009); Karlsson et al. (2024)). AIPW estimators model both the treatment-dependent conditional expected outcome and treatment probability and are doubly robust in the sense that if either of the estimators for these models is correctly specified, the AIPW estimator remains consistent. In our setting, where the design is randomized and the propensity score is known, a similar robustness guarantee holds. However, unlike the AIPW estimator, our approach relies on a prognostic cluster-level outcome model, $\mathbb{E}[\tilde{Y}\mid X]$, that does not incorporate treatment information, as would typically be required in a doubly robust estimator. The form of the augmented estimator also shares similarities with prediction-powered inference estimators (Angelopoulos et al., 2023), which use an auxiliary prediction model to improve inference, although those estimators do not employ any weighting function.
We can use any flexible, data-adaptive model to learn $\widehat{g}$, and notably, $\widehat{\phi}^{\text{augmented}}(\pi;\omega)$ is robust in the sense that it remains unbiased for the policy value regardless of whether $\widehat{g}(X)$ is correctly specified, provided that the weighting function $\omega$ satisfies some conditions. Additionally, to ensure this robustness guarantee, $\widehat{g}$ must be fitted independently of the observations used in the weighting estimator. In practice, this can be achieved using a cross-fitting procedure: the data are split into two folds, $\widehat{g}$ is fitted on one fold and predictions are made on the other fold, and then the roles of the folds are swapped and the procedure is repeated.
Theorem 4.
Suppose the weighted estimator $\widehat{\phi}(\pi;\omega)$ is unbiased for the policy value, $\mathbb{E}[\widehat{\phi}(\pi;\omega)]=V(\pi)$, and that the weighting function satisfies $\mathbb{E}[\omega(\boldsymbol{W},\boldsymbol{\pi},X)\mid X]=1$. Furthermore, assume that $\widehat{g}$ is obtained independently of the observations used to compute its predictions. Then the augmented weighting estimator $\widehat{\phi}^{\text{augmented}}(\pi;\omega)$ is unbiased for the policy value, $\mathbb{E}[\widehat{\phi}^{\text{augmented}}(\pi;\omega)]=V(\pi)$.
Proof.
We can show
$\mathbb{E}\left[\widehat{\phi}^{\text{augmented}}(\pi;\omega)\right]$ $=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\omega(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})(\tilde{Y}_{i}-\widehat{g}(X_{i}))+\widehat{g}(X_{i})\right]$
$=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\omega(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})\tilde{Y}_{i}\right]-\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\Big(1-\omega(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})\Big)\widehat{g}(X_{i})\right]$
$=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\omega(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})\tilde{Y}_{i}\right]-\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\Big[1-\omega(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})\mid X_{i}\Big]\widehat{g}(X_{i})\right]$
$=E\left[\frac{1}{N}\sum_{i=1}^{N}\omega(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})\tilde{Y}_{i}\right]$
$=E\Big[\widehat{\phi}(\pi;\omega)\Big]=V(\pi)$
where the fourth equality follows from the assumption that $\mathbb{E}\big[\omega(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})\mid X_{i}\big]=1$ and that $\widehat{g}$ can be taken out of the inner expectation because it is independent of the observations used to compute its predictions. The final equality follows from that the weighted estimator $\widehat{\phi}(\pi;\omega)$ is an unbiased estimator. ∎
To illustrate how the above theorem applies to the standard IPW and $\beta$-IPW estimators, we first note that the standard IPW estimator is unbiased under Assumption 1, and that $\mathbb{E}[\omega^{\text{IPW}}(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})\mid X_{i}]=1$ as shown in the proof of its unbiasedness in Appendix A.2. Under the additional Assumption 3 and the study design considered in this paper, the $\beta$-IPW estimator is also unbiased, and we can similarly show that $\mathbb{E}[\omega^{\beta\text{-IPW}}(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})\mid X_{i}]=1$ holds as follows:
$\mathbb{E}[\omega^{\beta\text{-IPW}}(\boldsymbol{W}_{i},\boldsymbol{\pi}_{i},X_{i})\mid X_{i}]$ $=\mathbb{E}\left[\sum_{\mathcal{U}\in\mathcal{I}_{i}^{\beta}}\prod_{j\in\mathcal{U}}\left(\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{\pi_{ij}}(X_{i})}-1\right)\mid X_{i}\right]$
$=\sum_{\mathcal{U}\in\mathcal{I}_{i}^{\beta}}\mathbb{E}\left[\prod_{j\in\mathcal{U}}\left(\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{\pi_{ij}}(X_{i})}-1\right)\mid X_{i}\right]$
[…]
$=1$
The third equality follows from explicitly excluding the empty set in the power set $\mathcal{I}i^{\beta}$ and noting that the product over the empty set is equal to one. The fifth equality follows from the fact that, conditional on $X_{i}$, the treatment WijW{ij} of unit $(i,j)$ is independent of all other treatments $W_{ij^{\prime}}$ for $j\neq j^{\prime}$. This independence allows the product to be moved outside the conditional expectation.
B.1 Experiment
Setup
We use the same data-generating process as in the main paper, with N=20, [^0] N=20{,}000 buyers (i.e., clusters) and $K=11$ items (i.e., units) per buyer. The interference structure is determined using the exponential decay function. To estimate $\mathbb{E}[\tilde{Y}\mid X]$, we use a linear logistic regression model, as the cluster-level outcome $\tilde{Y}$ is binary. We report the bias, variance, and mean-squared error averaged over 150 repetitions.
Results
As shown in Table 2, the augmented methods may reduce variance compared to their non-augmented counterparts. For the standard IPW estimator, augmentation leads to a decrease in variance (from 0.737 to 0.601) and a lower mean-squared error (from 0.738 to 0.603). For the $\beta$-IPW estimator, the variance reduction is negligible, likely because its variance is already low. Overall, augmentation with an outcome model appears to improve estimator efficiency, with the largest gains observed in cases where the initial variance is high, such as with the standard IPW estimator.
Table 2: Comparison of non-augmented and augmented estimators. We used $N=20.000$ number of buyers (i.e., clusters) and $K=11$ items (i.e., units) per buyer. We report the average bias, variance and mean-squared error with standard errors (in parentheses) over 150 repetitions.
Estimation strategy Bias Variance MSE
Standard IPW -0.003 (0.012) 0.737 (0.088) 0.738 (0.032)
Augmented Standard IPW 0.005 (0.010) 0.601 (0.068) 0.603 (0.022)
$\beta$-IPW ($\beta=1$) -0.016 (0.001) 0.005 (0.001) 0.005 (0.000)
Augmented $\beta$-IPW ($\beta=1$) -0.016 (0.001) 0.004 (0.001) 0.004 (0.000)
Appendix C Estimating Qini curves in settings with no interference
@@
Note that the above assumption is a special case of Assumption 2 with the exposure mapping $d_{ij}(\mathbf{W}_{i})=W_{ij}$.
@@
We can show that under Assumption 1 and 4, the above estimators are unbiased estimators for the policy value $V(\pi)$ and $C(\pi)$, respectively. Namely, we can show this with the same proof as in Appendix A.4, but replacing $d_{ij}(\mathbf{W_{i}})=[W_{ij},\overline{W}_{i}]$ with $d_{ij}(\mathbf{W_{i}})=W_{ij}$.
Appendix D Experimental details
D.1 Simulating marketplace dataset
@@
Next, to introduce heterogeneous treatment effects, we compute an item attractiveness score matrix $\mathbf{A}$ where element $A_{ij}$ relates buyer $i$’s interest in purchasing item $j$. This matrix depends on both the covariates and treatment as follows
$A_{ij}=\delta_{ij}\cdot(A_{ij}^{(0)}+W_{ij}\cdot A_{ij}^{(1)})~,$ (13)
where $A_{ij}^{(w)}=X_{i}^{\top}\Omega_{w}Z_{ij}$ with $\Omega_{w}\sim\text{U}([0,1]^{12\times 11})$ for $w\in\{0,1\}$. The variable $d_{ij}\sim\text{Bern}(0.5)$ randomly masks some elements in $A_{ij}^{(1)}$ to zero; this emulates that some items will not respond at all to a treatment. Here, $A_{ij}$ typically lies in the range $[0,1]$, but if necessary we clip it to this range so that we later could interpret it as a probability.
Simulating revenue and cost of treating items
In our simulations, we also compute the revenue and cost of items in the marketplace if they were to be treated. This allows us to evaluate treatment policies within the simulator that prioritize items based on their predicted expected profit. Specifically, we define the price and margin fraction of item $(i,j)$ as simple linear functions of its covariates $Z_{ij}$: the price is $P_{ij}=20+100\cdot Z_{ij,1}$ and the margin fraction is $M_{ij}=0.01+0.05\cdot Z_{ij,2}$, where $Z_{ij,1}$ and $Z_{ij,2}$ denote the first and second elements of $Z_{ij}$, respectively. If the treatment $W_{ij}=1$ corresponds to applying a fixed discount fraction of $d=0.08$, then the revenue of treating item $(i,j)$ is $R_{ij}=M_{ij}\cdot P_{ij}$ and the cost of treating item $(i,j)$ is $C_{ij}=d\cdot P_{ij}$. Consequently, the potential profit (or loss) from treating the item, conditional on it being converted, is $H_{ij}=R_{ij}-C_{ij}=(M_{ij}-d)\cdot P_{ij}$.
D.2 Treatment policy used for evaluation
To construct a simple baseline policy for our simulation studies, we define a treatment prioritization scoring rule $S_{\text{baseline}}$ that computes the expected profit from treating item $(i,j)$, ignoring the interference. Since our focus is on evaluating policies rather than developing them, this simplification avoids the complexity of implementing a policy that accounts for interference.
We then compute the scoring rule as
$S_{\text{baseline}}(X_{i},Z_{ij})=A_{ij}^{(1)}\cdot H_{ij}.$
Here, $A_{ij}^{(1)}$ represents the incremental change in the probability of conversion after an item has been treated (ignoring interference), while $H_{ij}$ denotes the profit from treating item $(i,j)$. The dependence on the covariates $(X_{i},Z_{ij})$ arises through both $A_{ij}^{(1)}$ and $H_{ij}$, as described in the previous subsection.
To generate treatment policies with varying performance, we introduce a parameter $\epsilon\in[0,1]$ used to perturb the baseline scoring rule. Specifically, we sample noise from a uniform distribution, $u\sim\mathcal{U}(S_{\text{min}},S_{\text{max}})$, where $S_{\text{min}}$ and $S_{\text{max}}$ are the minimum and maximum observed values respectively of the scoring rule $S_{\text{baseline}}(X_{i},Z_{ij})$. We then define a perturbed policy as
$S_{\text{baseline}}(X_{i},Z_{ij};\epsilon)=(1-\epsilon)\cdot S_{\text{baseline}}(X_{i},Z_{ij})+\epsilon\cdot u~.$
The best-performing policy corresponds to $S_{\text{baseline}}(X_{i},Z_{ij};\epsilon=0)$, while $S_{\text{baseline}}(X_{i},Z_{ij};\epsilon=1)$ is equivalent to a completely random policy.
D.3 Hardware used for experiments
All experiments have been performed on M1 Macbooks with 16GB of RAM. The runtime is highly dependent on the number of samples being simulated, but anything under 1 million samples takes not more than 10 minutes. The total runtime of our experiments, including all variants required for uncertainty estimation, took about 3 days worth of compute time.
D.4 Details on experiment shown in Figure 1
@@
We simulate a dataset with $N=20.000$ buyers (i.e., clusters) and $K=3$ items (i.e., units) per buyer with the exponential decay function. We estimate the Qini curve for the treatment prioritization rule $S_{\text{baseline}}(X_{i},Z_{ij};\epsilon=0)$ described in Appendix D.2 using the naive estimator and the standard IPW estimator. This was repeated 150 times and we plotted the average Qini curve.
D.5 Additional experimental results
In this section, we present additional experimental results to support the conclusions in Section 5. Figure 4 shows the variances of the estimated Qini curves as we vary either the number of buyers $N$ (i.e., clusters) or items $M$ (i.e., units). Meanwhile, Table 1 reports the average difference in AUC between the estimated Qini curves and the ground-truth curves over multiple repetitions. These results correspond to the setting with N=100, [^0] N=100{,}000 and $M=11$, which is the same as for the Qini curve plots shown in Figure 3 in the main paper.
@@ D.5 Additional experimental results
Figure 4: Comparing variance of each estimation strategy as we vary the number the buyers $N$ (i.e., clusters) or items $M$ (i.e., units) with $\eta_{\text{exp-decay}}$. While varying one, the other is kept fixed to either $N=20.000$ or $K=11$. The variance is reported over 150 repetitions.
Table 3: Comparison of estimation strategies via the Area Under the Curve (AUC) relative to the ground-truth Qini curve. Each entry reports the mean AUC and its standard error (in parentheses) across 150 repetitions. We let $N=100.000$ and $M=11$.
Interference structure Naive Standard IPW Fractional IPW $\beta$-IPW ($\beta=1$) $\beta$-IPW ($\beta=2$)
Max function 0.11 (0.02) 0.00 (0.05) 0.13 (0.04) -0.44 (0.02) -0.28 (0.02)
Product function 2.50 (0.03) -0.04 (0.07) 0.62 (0.06) -0.05 (0.02) -0.00 (0.03)
Exponential decay 0.17 (0.02) 0.01 (0.05) 0.11 (0.04) -0.18 (0.02) -0.07 (0.02)